what is the circumference of a circle with a diameter of 6cm?
Found 2 solutions by sarah_adam, checkley71:
Answer by sarah_adam(201)

Circumference of circle = 2*pi*r = d*pi (r = 2d)
C = (6)*(3.14) = 18.84 cm

Answer by checkley71(8403)

Circumferance of a circle is 2*pie*r or pie*d pie = 3.14 or 22/7, r=radius, d=diameter: Therefore for a diameter of 6cm diameter is circle=6*3.14 or
c=18.84cm